OpenStudy (anonymous):

Simplify only using positive exponents... (3^-5) * (3^-10) / 3^2

OpenStudy (anonymous):

\[3^{-5}\times 3^{-10}=3^{-5+-10}=3^{-15}\] and so \[\frac{3^{-15}}{x^2}=\frac{1}{x^{15}\times x^2}=\frac{1}{3^{17}}\]

OpenStudy (anonymous):

Hey, sorry to bother you satellite but I had one more question on my last question, could you help me with it?

OpenStudy (anonymous):

Once you get done.